# Imp: Highly Capable Large Multimodal Models for Mobile Devices

### v1

@@ 5 Main Results
The roadmap above ends with a combination of design choices that is generic and transferrable. In addition to Phi-2, we also apply these design choices to different lightweight LLMs, namely, Qwen-1.5 (1.8B) [2] and Phi-3 (3.8B) [47], to obtain a family of lightweight LMMs Imp-2B/3B/4B [^4] We have previously released an Imp model named Imp-v1-3B on HuggingFace (https://huggingface.co/MILVLG/imp-v1-3b), which uses slight different model configurations mentioned in the paper. To avoid ambiguity, the Imp models in this paper are termed Imp-v1.5-2B/3B/4B, respectively.. We conduct comprehensive quantitative and qualitative comparisons with state-of-the-art LMMs to validate the efficacy and efficiency of our Imp models. Each Imp model is trained on a server with 8 A100 GPUs (40GB) and finished in less than 32 hours.
@@ 5.1 Quantitative Comparisons with the State-of-the-art LMMs
Bunny-2B Qwen1.5-1.8B SigLIP-SO@384 2M 695K 76.6* 59.6* - - - - - 59.5 -

### v2

@@ 5 Main Results
The roadmap above ends with a combination of design choices that is generic and transferrable. In addition to Phi-2, we also apply these design choices to different lightweight LLMs, namely, Qwen-1.5 (1.8B) [2] and Phi-3 (3.8B) [47], to obtain a family of lightweight LMMs Imp-2B/3B/4B [^4] We have previously released an Imp model named Imp-v1-3B on HuggingFace (https://huggingface.co/MILVLG/imp-v1-3b). To avoid ambiguity, the Imp models obtained in this paper are termed Imp-v1.5-2B/3B/4B when released on HuggingFace. To be more precise, Imp-v1 and Imp-v1.5 use slightly different image preprocess strategies (“resize-to-square” for Imp-v1 and “resize-then-pad” for Imp-v1.5) and different training data (665K for Imp-v1 and 1M for Imp-v1.5). . We conduct comprehensive quantitative and qualitative comparisons with state-of-the-art LMMs to validate the efficacy and efficiency of our Imp models. Each Imp model is trained on a server with 8 A100 GPUs (40GB) and finished in less than 32 hours.
@@ 5.1 Quantitative Comparisons with the State-of-the-art LMMs
Bunny-2B Qwen1.5-1.8B SigLIP-SO@384 2M 695K 76.6* 59.6* - 64.6 - 85.8 1300.8 59.1 58.5 -